News by Date
Open and close menu tab
News by Category
Open and close menu tab
Press Resources
Open and close menu tab
Web Shortcuts
Open and close menu tab
Operations CXO Status Report

Friday 10 September 2004 7.30am EDT

During the last week Chandra completed the observing schedule as planned.

Real-time procedures were run on 8 Sept to perform a routine EIA Sequencer self-check and on 9 Sept as part of a diagnosis of a stuck-on heater located on the Optical Bench Assembly (OBA) in Zone 50.

A Chandra press release was issued on Sept 8 describing observations of core gas cloud motions in the Fornax galaxy cluster interpreted as attraction by a superstructure of dark matter. For details see: http://chandra.harvard.edu/press/04_releases/press_090804.html

The schedule of targets for the next week is shown below and includes two series of 21 calibration observations of AR Lac.

------------------------------------------
Swire Lockman 5        ACIS-I       Sep 12
Abell 478              ACIS-I       Sep 13
AR Lac (21 obs)        HRC-I
XRF040825              ACIS-I       Sep 14
Radiation Belts
Kepler                 ACIS-S       Sep 15
Sculptor Dwarf         ACIS-S
AR Lac (21 obs)        HRC-S
Swire Lockman 1        ACIS-I       Sep 16
Radiation Belts
Swire Lockman 2        ACIS-I       Sep 17
Swire Lockman 3        ACIS-I       Sep 18
CL0853+5759            ACIS-I       Sep 19

------------------------------------------

All spacecraft subsystems continued to support nominal operations.

Return to Status Reports